a right triangle has one leg of 29 length and one leg of 12 length, what is the angle measure of the angle farthest from the right angle?
Answer by Alan3354(69443)

a right triangle has one leg of 29 length and one leg of 12 length, what is the angle measure of the angle farthest from the right angle?
----------------
The right angle is C, the others are A & B
----
A = atan(12/29) =~ 22.479 degs
B = 90 - A =~ 67.52 degs